OpenStudy (anonymous):

Solve the following system using substitution or elimination: The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x. Be sure to show all work.

OpenStudy (anonymous):

We have perimeter is 46 so 2x+y=46 Also y is 4 more than x so y=x+4. Substitute this into the earlier equation to give 2x+(x+4)=46, i.e. 3x=42 so x=14, and then y =18

OpenStudy (anonymous):

thankyou so much!